1) Suppose that a charged particle of diameter 1 micron moving at constant speed in an electric field of magnitude 5 × 106 N/C experiences a drag force of magnitude 6 × 10-11 N opposing its motion. (1 micron = 10-6 m) What is the charge q on the particle? Ignore the effects of gravity.
Answer #1

Charged particle  of diameter 1 micron i.e we can treat it as a point charge.

electric field = 5 × 106 N/C

drag force of magnitude = 6 × 10-11 N (opposing its motion)

i.e the force is acting opposite of motion i.e the charge is negative

Magnitude Charge = Force / E =  6 × 10-11/  5 × 106 = 1.2*10-17 C

Hence Charge q = -  1.2*10-17 C